### v1

# Continual Learning for Multiple Modalities

Abstract
Continual learning aims to learn knowledge of tasks observed in sequential time steps while mitigating the forgetting of previously learned knowledge. Existing methods were proposed under the assumption of learning a single modality (e.g., image) over time, which limits their applicability in scenarios involving multiple modalities. In this work, we propose a novel continual learning framework that accommodates multiple modalities (image, video, audio, depth, and text). We train a model to align various modalities with text, leveraging its rich semantic information. However, this increases the risk of forgetting previously learned knowledge, exacerbated by the differing input traits of each task. To alleviate the overwriting of the previous knowledge of modalities, we propose a method for aggregating knowledge within and across modalities. The aggregated knowledge is obtained by assimilating new information through self-regularization within each modality and associating knowledge between modalities by prioritizing contributions from relevant modalities. Furthermore, we propose a strategy that re-aligns the embeddings of modalities to resolve biased alignment between modalities. We evaluate the proposed method in a wide range of continual learning scenarios using multiple datasets with different modalities. Extensive experiments demonstrate that ours outperforms existing methods in the scenarios, regardless of whether the identity of the modality is given.
1 Introduction
Recently, a vision-language model, pre-trained with a large number of image-text pairs, has drawn significant attention [28] due to its strong generalization performance [18, 29]. Leveraging the general knowledge of the pre-trained model, several algorithms have achieved promising results in image classification [14], semantic segmentation [21], and video action recognition [29], to name a few. However, when deep learning models are exposed to new data, they tend to overwrite previously learned knowledge, leading to forgetting earlier information. This issue is known as catastrophic forgetting [22, 16].
Figure 1: Illustration of (a) existing unimodal and (b) the proposed multimodal continual learning. Multimodal continual learning poses a significant challenge, as it complicates the retention of previously acquired knowledge while integrating information from tasks across diverse modalities.
To alleviate this phenomenon, continual learning methods strive to preserve existing knowledge while accommodating new information [16, 1, 30, 33, 11, 17]. To this end, they apply penalties to minimize changes in key parameters that contribute to performance on previous tasks [16, 1] or distill the knowledge of the old model into newer ones [11, 17]. Furthermore, to facilitate recall of previous knowledge, some methods rely on having access to data from old tasks [23, 30, 7, 20]. The collected data is then used jointly with new data to train the model. However, storing raw data can lead to privacy concerns in many applications [34]. Moreover, temporal data (e.g., video and audio) demand more memory resources than non-temporal data [37, 13], posing additional challenges for continual learning.
Recently, continual learning methods that adapt a pre-trained model (e.g., ViT [6] or CLIP [28]) have been proposed [44, 43, 40]. These methods are designed to select prompts from a shared pool for each task. The selected prompts are inserted into a pre-trained model and updated to acquire new knowledge. Despite outperforming other methods by leveraging the general knowledge of the pre-trained model, these approaches validate on a single modality and do not account for data from new modalities. This limitation restricts their applicability in dynamic environments where data from different modalities appear, such as in autonomous driving [26] and robot planning [31]. Furthermore, these approaches are susceptible to forgetting knowledge learned previously when learning multiple modalities because modality gaps between tasks cause prompts trained on different input types to interfere with each other.
In this paper, we propose a novel approach, COMM (COntinual Learning for Multiple Modalities), for continuously learning tasks of different modalities, using a vision-language model. Specifically, we take text-bound data as input, leveraging the rich semantic information provided by text [53]. We process the text data through a language encoder and the other modalities through a modality encoder, enabling the integration of additional modalities without introducing many parameters. However, directly updating the parameters of the encoder across modalities can lead to interference due to differing input traits. To address this issue, we propose a prompt aggregation strategy that associates knowledge within modality through self-regularization, gradually acquiring newly introduced knowledge. It then integrates relevant knowledge between modalities while minimizing the influence of unrelated ones. We introduce a projection head for each modality to assist in aligning embeddings from both encoders, following the practice in [8, 9]. Additionally, to avoid the biased projection of modality data towards text data learned most recently [46], we present a projection head alignment strategy that re-aligns previous modality embeddings with their corresponding text embeddings. Figure 1 illustrates the proposed multimodal continual learning compared to existing unimodal continual learning.
We demonstrate the proposed method by comparing it with existing continual learning methods across tasks involving multiple modalities using classification benchmark datasets: ImageNet-100 [5], UCF-101 [36], SUN-RGBD [35], and ESC-50 [25]. Experimental results show that the proposed method outperforms its competitors across various class-incremental learning scenarios while introducing negligible memory overhead from the prompts. The contributions of our work are four-fold:
  • We propose the first multimodal continual learning method that leverages multiple modalities, addressing the challenge of forgetting that arises from learning the distinct traits of each modality in sequential data.
  • We present a knowledge aggregation approach that effectively learns a new task by mitigating interference between modalities while preserving existing knowledge.
  • We self-regulate prompts to consolidate them and selectively prioritize those from relevant modalities, enabling optimal use of prompts within and across modalities.
  • Extensive experiments show that the proposed method achieves outstanding performance compared to existing continual learning methods by effectively preventing detrimental knowledge mixing across modalities.
2 Related Work
Multimodal learning. Multimodal learning aims to align representations of multiple modalities [51, 14, 29, 19, 45, 10]. They have recently gained attention with the emergence of a model pre-trained on image-text pairs [28]. Several methods [14, 51, 52, 45] have been proposed to adapt the pre-trained model to downstream image classification tasks. Additionally, methods [19, 29] have been proposed to adapt image-text matching models to facilitate video-text matching. They fine-tune pre-trained weights or insert learnable prompts to adapt to downstream tasks. Some recent works [10, 9, 53] have proposed learning multiple modalities (three or more, including image and text). These methods integrate representations extracted using multiple networks, each tailored to a specific modality. Although the methods train modalities simultaneously, training them sequentially presents a greater challenge, as it requires maintaining consistent representations for new modality data while preserving those learned from previously encountered modalities. Additionally, employing a modality-specific network limits scalability as the number of modalities increases. In this work, we present a global encoder for non-text modalities that aligns prompts from relevant modalities, reducing interferences and promoting consensus across different modalities.
Continual learning. The goal of continual learning is to maintain previously learned knowledge while simultaneously acquiring new information. Most existing continual learning methods were proposed based on a single modality (e.g., image). Early works proposed to identify contributing parameters for previous tasks [16, 48, 2, 1]. These methods train the model by penalizing updates to key parameters, helping to alleviate the forgetting of previously learned knowledge. However, they can result in significant loss of knowledge as the number of tasks increases when the semantic domains of the tasks differ [4]. Several studies store [7, 30, 20, 24] or generate [23, 32, 42] some data from old tasks and train the network with a joint set of old and new data. However, privacy or memory budget issues can arise from storing old data [37]. Additionally, these approaches learn a single modality, which limits their applicability in real-world data involving multiple modalities.
Continual learning with a vision-language model. Recent approaches have been proposed using a pre-trained vision-language model [28] to classify unimodal data by matching it with text [50, 45, 38, 40, 27]. Some methods [50, 45] directly update the parameters of a model while regularizing to retain the learned information of the model. The others [38, 40, 27] learn new tasks by combining trainable prompts with the input embeddings in each layer of a pre-trained model, selecting prompts from a predefined pool. However, these methods overwrite previously learned knowledge by updating the parameters of the model (or previously learned prompts), thereby exacerbating forgetting due to interference from different modalities across tasks. We aggregate prompts from modalities relevant to the input by ensuring that knowledge learned from other modalities remains unchanged, alleviating interference between modalities.
3 Methodology
Figure 2: An illustration of the proposed method for learning multiple modalities ($D^{t}_{\color[rgb]{0,1,1}\textbf{depth}},D^{t}_{\color[rgb]{0,1,0}\textbf{video}},...$) at time step $t$. We introduce a prompt for each modality, which is aggregated with existing prompts, and perform self-regulation to preserve previous knowledge of modalities. To prioritize the knowledge of relevant modalities, we associate the aggregated prompts across modalities. The data and the aggregated prompts are fed into the modality and language encoders to obtain modality and text embeddings, respectively. We also present projection head re-alignment to alleviate biased projections toward the most recently learned text.
3.1 Problem Definition
[…]
Modality-text alignment. To extract embeddings of the modality data $x^{t}_{m,i}$ and the text data $c^{t}_{text,j}$, we use the modality [^1] We implement the modality encoder using the vision encoder in [28]. and language encoders, $V(\cdot)$ and $L(\cdot)$, in a pre-trained model [28], respectively. This approach helps align embeddings from non-text modalities with those from text, following [9, 53]. To mitigate the increase in memory cost as the number of modalities grows, non-text modalities share a common modality encoder. $x^{t}_{m,i}$ and $c^{t}_{text,j}$ are preprocessed and tokenized as $e^{t}_{m,i}$ and $w^{t}_{text,j}$. By using the tokenized data, we obtain the modality feature $\bar{v}^{t}_{m,i}=V(e^{t}_{m,i})$ and the text embedding $l^{t}_{text,j}=L(w^{t}_{text,j})$, respectively. $\bar{v}^{t}_{m,i}$ is further projected to a common modal-text embedding space as $v^{t}_{m,i}=f_{\alpha^{t}_{m}}(\bar{v}^{t}_{m,i})$, where $f_{\alpha^{t}_{m}}(\cdot)$ is a projection head parameterized by $\alpha^{t}_{m}$. To learn the modality data, we maximize the predictive probability defined as
$p(y^{t}_{i}|x^{t}_{m,i})=\frac{\mbox{exp}(\mbox{sim}(v^{t}_{m,i},l^{t}_{text,y^{t}_{i}})/\tau)}{\sum_{j}\mbox{exp}(\mbox{sim}(v^{t}_{m,i},l^{t}_{text,j})/\tau)},$ (1)
where $l^{t}_{text,y^{t}_{i}}$ denotes the text embedding of the label $y^{t}_{i}$. $\mbox{sim}(\cdot)$ and $\tau$ represent the cosine similarity and the temperature parameter of the pre-trained model, respectively.
3.2 Framework
The proposed method learns multimodal tasks based on a pre-trained model. However, directly updating the model on different modalities [16] can exacerbate forgetting due to modality gaps and differing input sizes across tasks. To resolve it, we take a strategy of introducing prompts, learnable parameters, to learn new tasks without updating the model [14]. Figure 2 illustrates the proposed continual learning method for multiple modalities.
Specifically, the set of modalities of the task [^2] We assume that a task appears at every time step in continual learning. at time $t$ is $\mathcal{M}^{t}$, where each modality consists of text-paired data. To learn new tasks, we introduce and update a pair of learnable prompts, $P^{t}_{m}$ and $Q^{t}_{m}$, of the modality and language encoders by incorporating them with the input tokens, respectively. Unlike existing prompt-based continual learning [52, 40], which employs a shared prompt pool for unimodal tasks, multimodal continual learning requires a shared pool that can accommodate tasks with diverse modalities. This introduces a challenge in selecting appropriate prompts.
To effectively utilize prompts for each modality and minimize interference with previously learned representations, we propose aggregating prompts within and between modalities in the proposed pipeline. In the first stage, to alleviate significant changes in old knowledge across time within a modality, we combine the prompts learned so far with newly introduced ones by a self-regulating strategy. Note that updating prompts across different modalities can significantly alter the knowledge of the prompts due to gaps in input characteristics, making it more susceptible to forgetting. To address this, in the second stage, we associate knowledge between modalities by leveraging relevant modality connections. We obtain modality and text embeddings using prompts aggregated within and between modalities. The embeddings are aligned by updating $P^{t}_{m}$ and $Q^{t}_{m}$ using the cross-entropy loss,
$\mathcal{L}_{m}=-\sum_{i}y^{t}_{i}\log p(y^{t}_{i}|x^{t}_{m,i}).$ (2)
3.3 Knowledge Aggregation
The use of prompts from a shared prompt pool [44, 40] over time presents two challenges: 1) difficulty in learning new tasks due to the inadequate utilization of existing prompts, and 2) performance degradation as prompts learned for earlier tasks are updated. To address these issues, we propose a knowledge aggregation strategy that minimizes interference for each modality and across modalities over time.
Aggregation within a modality. To maintain earlier representations and ensure effective prompt utilization within a modality, we aggregate prompts over time and self-regulate them. Specifically, prompts $P^{t}_{m}$ and $Q^{t}_{m}$ are zero-initialized to gradually acquire knowledge of the task. These prompts are then integrated with those from the previous tasks as ¯$\bm{\bar{}}{P}^{t}_{m}=\oplus(\bm{\bar{}}{P}^{t-1}_{m},P^{t}_{m})$ and ¯$\bm{\bar{}}{Q}^{t}_{m}=\oplus(\bm{\bar{}}{Q}^{t-1}_{m},Q^{t}_{m}$), where ¯$\bm{\bar{}}{P}^{1}_{m}=P^{1}_{m}$ and ¯$\bm{\bar{}}{Q}^{1}_{m}=Q^{1}_{m}$. $\oplus$ denotes element-wise summation. Prompts that were not introduced during previous time steps are excluded from the summation. The modality-text embeddings, $V^{t}_{i,\bm{\bar{}}{P}^{t^{\prime}}_{m}}$triangleq $f_{\alpha^{t}_{m}}(V(e^{t}_{m,i};\bm{\bar{}}{P}^{t^{\prime}}_{m}))$ and $L^{t}_{i,\bm{\bar{}}{Q}^{t^{\prime}}_{m}}$triangleq $L(w^{t}_{text,y^{t}_{i}};\bm{\bar{}}{Q}^{t^{\prime}}_{m})$, are aligned by minimizing the loss, $\mathcal{L}_{m}$. However, updating the prompts by the loss does not guarantee the preservation of previous representations. To closely align the integrated knowledge with the previous knowledge, we update prompts with self-regularization as
$\mathcal{L}_{sr}=\sum_{i}||V^{t}_{i,\bm{\bar{}}{P}^{\bm{t-1}}_{m}}-V^{t}_{i,\bm{\bar{}}{P}^{\bm{t}}_{m}}||_{2}+||L^{t}_{i,\bm{\bar{}}{Q}^{\bm{t-1}}_{m}}-L^{t}_{i,\bm{\bar{}}{Q}^{\bm{t}}_{m}}||_{2}.$ (3)
By incorporating the self-regularizations into $\mathcal{L}_{m}$, we derive a representative prompt $\bar{P}^{t}_{m}$ for each modality $m$ that is free from incorrect prompt selection within the modality. Similarly, we aggregate old and new projection heads to form ¯$\bm{\bar{}}{\alpha}^{t}_{m}$= (¯$\oplus(\bm{\bar{}}{\alpha}^{t-1}_{m},\alpha^{t}_{m})$ and apply self-regularization to minimize the distance between outputs extracted using ¯$\bm{\bar{}}{\alpha}^{t}_{m}$ and ¯$\bm{\bar{}}{\alpha}^{t-1}_{m}$ for the input $V(e^{t}_{m,i};\bm{\bar{}}{P}^{t}_{m})$. This approach ensures that both the prompts and projection heads evolve gradually, retaining previously learned knowledge within the modality.
Aggregation between modalities. Prompts trained across modalities are more sensitive to inappropriate usage than those trained for a single modality. This necessitates the identification of relevant modalities from the pool observed thus far, denoted as $\mathcal{M}^{1:t}$. To do so, we introduce a function $g_{\beta^{t}}(\cdot)$: $\mathbb{R}^{d_{i}}\rightarrow\mathbb{R}^{|\mathcal{M}^{1:t}|}$, parameterized with $\beta^{t}$, to identify relevant modalities by examining the features extracted from the modality encoder, $V(\cdot)$. The function $g_{\beta^{t}}(\cdot)$ assigns probabilities to the prompts, representing their relevance and facilitating the integration of prompts from the modalities. Due to the inaccessibility of old data, we collect features for each modality, which are sampled [49] from a normal distribution defined by the mean and covariance of the previously observed features, to update $g_{\beta^{t}}(\cdot)$
$\mathcal{L}_{agg}=\sum_{m^{\prime}\in\mathcal{M}^{1:t}}\sum_{i=1}^{\bar{n}_{m^{\prime}}}~{}-\log g_{\beta^{t}}(m^{\prime}|\bar{u}_{m^{\prime},i}),$ (4)
where $\bar{u}_{m^{\prime},i}$ is a collected sample and $\bar{n}_{m^{\prime}}$ denotes the number of samples. By minimizing this loss, $g_{\beta^{t}}$ yields relevance of each modality $m^{\prime}\in\mathcal{M}^{1:t}$ in relation to $m$. Then we aggregate the prompts across modalities with the relevance $g_{\beta^{t}}$ as:
^$\bm{\hat{}}{P}^{t}_{m}=\sum_{m^{\prime}\in\mathcal{M}^{1:t}}g_{\beta^{t}}(m^{\prime}|V(e^{t}_{m,i}))\cdot\bm{\bar{}}{P}^{t}_{m^{\prime}},$ (5)
which guides the contribution of each prompt to the aggregated prompt ^$\bm{\hat{}}{P}^{t}_{m}$. Note that the text prompts do not require aggregation across modalities because text data is input solely into the language encoder. By aligning the modality and text embeddings extracted with the prompts, ^$\bm{\hat{}}{P}^{t}_{m}$ and ¯$\bm{\bar{}}{Q}^{t}_{m}$, we can avoid interference between modalities. To summarize, we update the set of parameters $\{P^{t}_{m},Q^{t}_{m},\alpha^{t}_{m}\}$ by minimizing the aforementioned loss functions.
Re-alignment modality projection head. Learning new knowledge without accessing data from old tasks can cause the embeddings from the modality encoder to be biased toward the text embeddings of the latest task. To address this issue, we introduce an additional phase to re-align the projection head $f_{\bm{\bar{}}{\alpha}^{t}_{m}}(\cdot)$ so that old and new modality embeddings are distinguished. Specifically, for each class of modality $m$, we collect features mimicking old features, which are sampled using $V(\cdot;\bm{\hat{}}{P}^{i}_{m})$, where $i<t$. The aggregated projection head, ¯$\bm{\bar{}}{\alpha}^{t}_{m}$, is further fine-tuned to align the modality embedding derived from the sampled features with the text embedding by minimizing $\mathcal{L}_{m}$.
4 Experiments
4.1 Setup
Datasets. We evaluated the proposed method for continual learning of five modalities: image, video, audio, depth, and text. The datasets used for the modalities are ImageNet-100 [5], UCF-101 [36], ESC-50 [25], and SUN-RGBD [35], each accompanied by corresponding text data. ImageNet-100 contains 100 classes selected from the original ImageNet dataset [5], widely used in continual learning for image classification [7, 47]. UCF-101 contains 101 action categories for video action recognition [36]. ESC-50 is a fine-grained collection of environmental audio recordings of 50 classes for sound classification. For the depth modality, we performed on the SUN-RGBD dataset for scene classification. Each dataset was divided into five non-overlapping subsets of classes. We consider each subset as a task. Please refer to the supplementary material for preprocessing each modality dataset and the class orders.
Scenarios. We compared the proposed method with other methods with three sequences of tasks: (i) A random modality is introduced at each time step. (ii) A modality is trained across multiple time steps before shifting to the next modality. (iii) All modalities are available at every time step. The first two sequences comprise 20 time steps (corresponding to five subsets of four modalities), with each task exclusively learning a single modality. In contrast, the third sequence consists of five time steps, during which all modalities are learned simultaneously for every step. For each sequence, we evaluated the methods with and without the provision of modality identity during the evaluation phase, which we refer to as the modality-specific and modality-agnostic scenarios, respectively. We report average incremental accuracy (AIA) [30, 7], final average accuracy (FAA) [7, 39], and average forgetting (F) [3, 33].
Implementation details. For a fair comparison with other continual learning methods, we adopted ViT-B-16 CLIP [28] as the backbone model for the compared methods. We used the vision and text encoders of CLIP for the non-text and text modalities, respectively, in accordance with established practices [9, 53]. We used a fully connected layer for $g_{\beta^{t}}(\cdot)$. Additionally, due to the high dimension of the projection layer $\alpha^{t}_{m}$, we employed a low-rank decomposition technique [12] to split it into two smaller matrices with a rank of 20. The proposed method was trained using the Adam optimizer [15] with $\beta_{1}$ and $\beta_{2}$ of 0.9 and 0.999. We report the average results from three independent runs for all experiments using NVIDIA RTX A5000 GPUs. For additional details on the implementation, including the preprocessing strategies, please refer to the supplementary material.
Baselines. We evaluated the proposed method, COMM, with existing continual learning methods, including sequential fine-tuning (FT), EWC [16], LwF [17], WISE-FT [45], L2P [44], S-liPrompts [41], AttriCLIP [40], and PGP [27]. For the conventional methods, EWC, LwF, and WISE-FT, we directly updated the parameters of modality and text encoders. The prompt-based methods, L2P [^3] Since L2P is a ViT-based method [6], we incorporated the prompts into the modality encoder in the proposed framework for comparison., AttriCLIP, and PGP, used modality-specific and modality-shared prompt pools for the modality-specific and modality-agnostic class-incremental learning scenarios.
Table 1: Results for the first sequence on the modality-specific (top) and modality-agnostic (bottom) class-incremental learning scenarios. The AIA, FAA, and F metrics are measured after training all tasks. Overall denotes the average AIA and FAA across all modalities.
Method Image Video Depth Audio Overall
AIA ($\uparrow$) FAA ($\uparrow$) F ($\downarrow$) AIA ($\uparrow$) FAA ($\uparrow$) F ($\downarrow$) AIA ($\uparrow$) FAA ($\uparrow$) F ($\downarrow$) AIA ($\uparrow$) FAA ($\uparrow$) F ($\downarrow$) AIA ($\uparrow$) FAA ($\uparrow$) F ($\downarrow$)
Modality-Specific Class-incremental Learning
FT 45.66 23.46 23.37 45.25 13.16 33.87 26.21 7.58 19.72 28.96 4.20 26.31 36.52 12.16 25.82
EWC [16] 47.16 23.42 24.98 48.59 20.76 29.37 26.38 13.91 13.20 26.68 5.50 20.50 37.20 15.90 22.51
LwF [17] 53.13 24.78 29.84 48.59 23.26 26.74 30.64 11.43 20.34 31.89 9.75 23.53 41.06 17.31 25.11
WISE-FT [45] 46.73 21.96 26.08 46.99 21.27 27.14 26.93 5.27 22.93 22.69 8.25 15.34 35.83 14.18 22.87
L2P [44] 76.96 68.66 8.74 78.49 70.71 8.21 45.06 24.75 24.66 44.28 17.00 28.98 61.20 45.28 17.65
S-liPrompts [41] 75.61 65.25 10.91 82.85 69.04 14.57 47.98 26.08 30.07 42.68 22.25 21.70 62.28 45.66 19.31
AttriCLIP [40] 77.39 65.70 12.30 77.40 70.51 7.28 51.21 35.01 17.16 43.44 20.50 24.37 62.36 47.93 15.27
PGP [27] 78.88 68.58 10.84 83.14 71.46 12.33 48.67 25.58 24.45 47.69 26.25 22.78 64.60 47.98 17.60
COMM (Ours) 86.07 78.96 7.11 91.03 83.82 7.21 62.90 44.03 18.88 61.86 43.25 18.60 75.47 62.51 13.71
Modality-Agnostic Class-incremental Learning
FT 33.39 1.46 33.61 27.90 0.14 29.30 25.70 7.59 19.18 23.65 4.25 20.61 27.66 3.36 25.68
EWC [16] 38.23 5.60 34.34 27.46 3.63 25.15 24.04 13.91 10.73 22.77 5.50 18.34 28.13 7.15 22.14
LwF [17] 48.74 21.14 29.05 19.82 1.17 22.43 23.01 6.74 17.22 19.01 9.00 10.70 27.65 9.51 19.85
WISE-FT [45] 38.43 10.34 29.57 24.89 0.36 25.88 25.94 5.27 21.67 22.46 6.25 17.22 27.93 5.55 23.59
L2P [44] 79.85 69.84 10.53 71.25 65.14 6.44 39.92 26.72 13.96 33.08 21.50 12.31 56.03 45.80 19.31
S-liPrompts [41] 77.36 66.72 11.20 79.52 73.18 6.69 40.94 30.02 11.56 33.63 23.50 10.79 57.87 48.36 10.06
AttriCLIP [40] 73.93 63.90 10.56 73.28 68.70 4.83 39.52 28.57 11.60 34.99 22.50 13.27 55.43 45.91 10.06
PGP [27] 81.29 70.92 10.85 77.55 67.18 10.94 45.61 26.25 20.49 46.15 21.50 26.19 62.37 41.58 17.12
COMM (Ours) 83.89 76.84 5.43 90.83 83.63 7.60 62.78 43.94 19.84 61.85 43.25 19.60 74.84 61.92 13.19
Figure 3: Results of modality-agnostic continual learning for the first sequence, where FAA is measured at each time step. The vertical lines represent the time steps at which the modality is learned.
4.2 Main Results
Sequence 1. We evaluated our method and the compared methods on the first sequence for multimodal continual learning. Table 1 summarizes the results of the modality-specific (top) and modality-agnostic (bottom) class-incremental learning scenarios. Overall, we observe that the results of the modality-agnostic scenario show lower performance compared to the modality-specific scenario due to the challenge in identifying the modality. The regularization methods, EWC, LwF, and WISE-FT, perform poorly in both scenarios compared to the prompt-based continual learning methods, L2P, S-liPrompts, AttriCLIP, PGP, and COMM. This indicates that directly updating the network parameters in the regularization methods to accommodate knowledge from multiple modalities is prone to forgetting previously learned tasks due to contaminated mixed knowledge. Even though the existing prompt-based methods outperform other continual learning approaches, they do not perform well for the depth and audio modalities. Notably, these methods exhibit lower AIA in the modality-agnostic scenario due to the challenge of selecting prompts from the shared pool, which hinders learning new tasks. The proposed method, aiming to alleviate sensitivity to prompt selection and biased projections, outperforms the best competitor, S-liPrompts, with an overall FAA gap of 16.85% and 13.56% in the respective scenarios. In terms of the performance gap between the scenarios, the proposed method shows only a marginal performance gap of 0.63% in overall AIA, in contrast to the larger gaps of 4.41% and 6.93% observed with S-liPrompts and AttriCLIP, respectively. Additional results for different class orders can be seen in the supplementary material.
We also present detailed results of the modality-agnostic scenario (measured in FAA) of the compared methods in Figure 3. The conventional methods that directly update the network parameters exhibit high fluctuations. The performance of each modality slightly increases when tasks with the specific modality are learned (indicated by vertical lines). However, learning tasks from other modalities (indicated by non-vertical lines) results in a significant performance decline due to overwritten knowledge. The prompt-based continual learning methods, L2P, S-liPrompts, AttriCLIP, and PGP, also show unsatisfying performance in video, depth, and audio modalities. In contrast, the proposed method exhibits stable performance across different modalities, surpassing other approaches, particularly in the audio and depth modalities.
Table 2: Results for the second sequence. AIA and F are measured after training on all tasks. Overall AIA and Overall F represent the average AIA and average F across all modalities, respectively.
Method Modality-Specific Learning Modality-Agnostic Learning
Image ($\uparrow$) Video ($\uparrow$) Depth ($\uparrow$) Audio ($\uparrow$)
Overall
AIA ($\uparrow$)
Overall
F ($\downarrow$)
Image ($\uparrow$) Video ($\uparrow$) Depth ($\uparrow$) Audio ($\uparrow$)
Overall
AIA ($\uparrow$)
Overall
F ($\downarrow$)
FT 26.38 27.27 20.00 26.39 25.01 24.56 26.05 22.61 16.45 25.74 22.71 22.64
EWC [16] 27.21 28.45 18.19 30.32 26.04 24.86 26.10 24.37 16.98 29.88 24.33 26.59
LwF [17] 26.92 28.52 18.71 27.76 25.48 24.83 20.48 23.12 18.23 24.45 21.57 21.49
WISE-FT [45] 27.08 27.85 16.53 29.17 25.16 24.67 23.43 24.00 16.01 28.02 22.87 24.63
L2P [44] 70.73 72.96 32.25 39.54 53.87 11.25 72.19 62.37 37.03 43.38 53.74 11.95
S-liPrompts [41] 67.93 75.45 40.88 43.25 56.88 12.02 67.81 73.42 32.17 38.62 53.01 11.29
AttriCLIP [40] 68.62 72.69 42.30 42.24 56.46 10.43 67.89 69.91 38.26 44.05 55.03 12.62
PGP [27] 72.51 76.51 37.42 43.87 57.58 10.11 72.15 66.24 37.22 42.72 54.58 11.31
COMM (Ours) 81.95 85.17 50.39 61.01 69.63 7.81 80.35 84.99 50.30 60.94 69.14 7.94
Figure 4: Results (measured in FAA) of multimodal continual learning for the third sequence.
Sequence 2. To further evaluate the consistency of our approach, we compared COMM with other continual learning methods in another sequence, where each modality is trained over multiple times before shifting to another (in order of image, video, depth, and audio). Table 2 reports the experimental results for the sequence, demonstrating a similar performance trend to that observed in Table 1. However, conventional continual learning methods show a significant performance degradation for modalities learned early. Specifically, the image and video modalities experience about a 20% drop in AIA compared to the first sequence. Similarly, prompt-based methods show performance declines for previously learned modalities; however, they achieve enhanced performance for the most recently acquired modality, audio. The results indicate that learning different modalities increases interference between modalities, leading to more forgetting in both continual learning methods. COMM outperforms PGP and AttriCLIP in each scenario, demonstrating significant AIA improvements in audio with gaps of 17.14% and 16.89%, respectively. Additionally, it achieves overall AIA improvements of 12.05% (with an overall F decrease of 2.30%) and 14.11% (with an overall F decrease of 4.68%) compared to PGP and AttriCLIP. We also provide additional results on sequences with varying task orders in the supplementary material.
Table 3: Results of the ablation study on the usage of the components in the proposed method.
Method First Sequence (Random) Second Sequence (Shift)
AGG SR PHR Image Video Depth Audio Overall Image Video Depth Audio Overall
- - 62.81 80.42 46.44 49.84 59.88 59.77 70.14 23.35 43.76 49.26
- - 73.95 75.14 48.92 59.35 64.34 65.52 65.94 30.75 56.24 54.61
- - 76.10 81.50 49.32 53.65 65.14 67.81 76.88 31.92 51.11 56.93
- 71.91 73.21 48.51 59.53 63.29 64.08 65.78 41.76 54.46 56.53
- 76.40 81.47 58.01 54.99 67.72 66.73 76.63 25.65 49.77 54.70
- 83.05 88.62 47.55 60.59 69.95 78.17 81.51 29.92 60.83 62.61
83.89 90.83 62.78 61.85 74.84 80.35 84.99 50.30 60.94 69.14
Table 4: Results of the ablation study on prompting designs.
Method Image Video Depth Audio Overall
COMM-L 85.30 83.89 61.56 60.43 72.80
COMM-V 85.76 90.50 59.22 61.35 74.21
COMM-VL 86.07 91.03 62.90 61.86 75.47
Sequence 3. We report the results for multiple modalities learned simultaneously at every step in a modality-specific scenario. Figure 4 shows the results for the sequence. The results reveal a clear difference between methods using prompts (L2P, S-liPrompts, AttriCLIP, and PGP) and those updating model parameters (EWC, LwF, and WISE-FT) across all time steps. This indicates that directly updating model parameters for multiple modalities can lead to severe conflicts, resulting in poor performance. The proposed approach consistently demonstrates higher performance across all modalities than the competitors due to reduced interference between modalities.
4.3 Ablation Study and Analysis
Ablation study on the components. We conducted an ablation study of the proposed method, including its components: prompt aggregation (AGG), self-regularization (SR), and projection head re-alignment (PHR). Note that the method without AGG or SR requires identifying the modality identity or task to select prompts, respectively. Therefore, we employed a prompt selection mechanism in [41]. We compared these methods to the first two sequences in the modality-agnostic scenario. Table 3 reports the results of the ablation study with AIA measured after training all tasks. Overall, the proposed method using an individual component results in unsatisfactory performance on both task sequences compared to employing them together. The method using AGG identifies relevant modalities but struggles to select accurate task-specific prompts. When comparing the method utilizing all components against that excluding AGG, it becomes evident that selecting suitable modality prompts significantly enhances performance, with gaps of 15.23% and 20.38% in the depth modality for the first and second sequences, respectively. The method that employs both SR and PHR exhibits a strong synergistic effect because SR preserves previous features by regularizing prompts, and these preserved features closely resemble those used by PHR. We provide visualizations demonstrating the maintainability of SR in the embedding space in the supplementary material.
Figure 5: Results for the proposed method with varying learnable parameters by adjusting the prompt dimension (top) and the layers where prompts are inserted (bottom).
Ablation study on the prompting designs. We conducted an additional study to investigate the impact of prompting design strategies in continual learning. In this study, we compared the proposed method, which uses prompts for both modality and language encoders (which we call COMM-VL), with the method that uses prompts for the modality (COMM-V) or the language encoder (COMM-L). We applied these methods to the first sequence in the modality-specific scenario. In Table 4, we observe that COMM-VL yields higher overall AIA performance than COMM-V and COMM-L. COMM-L gives the lowest performance among the compared approaches due to its insufficiency in learning diverse modalities, highlighting the requirements for prompts in the modality encoder.
Study of scalability. In this study, we evaluated the scalability of the proposed method by examining the impact of varying numbers of learnable parameters. To adjust the number of learnable parameters, we increased the dimension of prompts and the layers of the pre-trained models for inserting prompts, following the methods in the previous studies [14, 51]. Figure 5 indicates that COMM maintains consistent overall AIA performance across different prompt dimensions and numbers of layers for prompt insertion. Furthermore, we also compared the parameter growth of COMM with other prompt-based methods in Figure 6. Overall, prompt-based methods require fewer parameters than updating the entire model [16, 17], which involves approximately 124 million parameters. Specifically, L2P and AttriCLIP leverage a prompt pool for each modality, resulting in a step-wise increase in parameters as new modalities are introduced. S-liPrompts uses task-specific prompts, leading to a monotonic parameter increase. The proposed method, COMM, consolidates prompts inter- and intra-modalities, enhancing scalability when integrating additional modalities.
Figure 6: The number of parameters with respect to the modality.
5 Conclusion
In this work, we have introduced the first multimodal continual learning framework that addresses multiple tasks of different modalities. Learning multiple tasks complicates the retention of prior knowledge as accumulating mixed information over time interferes with the learning of individual modalities. To address this challenge, we have proposed a strategy that aggregates knowledge in each modality over time through self-regularization and integrates it across relevant modalities. This approach selectively emphasizes contributions from relevant modalities while minimizing the influence of unrelated ones, gradually assimilating new information and preserving existing knowledge. Extensive experimental results, including modality-specific and modality-agnostic scenarios, demonstrated that ours significantly outperforms other continual learning methods with a notable performance gap.
Supplementary Material
A Additional Implementation Details
A.1 Preprocessing multiple modalities
Images from ImageNet-100 were resized to a spatial dimension of 224$\times$224 pixels. We applied center cropping for augmentation. For SUN-RGBD, we converted the depth maps into disparity maps following [8] and normalized the disparity map between 0 and 1. For the video dataset, we randomly selected three frames from every video clip of two seconds. The height and width of each video frame were resized to $224\times 224$ pixels, resulting in dimensions of $6\times 3\times 224\times 224$. We augmented videos with center cropping and horizontal flipping following [37]. For the audio dataset, we selected three audio clips from each audio sample. We processed each audio clip by sampling it at 16kHz, followed by extracting a mel-spectrogram with 224 frequency bins using a 25ms Hamming window every 10ms [9], and resized the spatial dimension to $224\times 224$, resulting in dimensions of $3\times 3\times 224\times 224$. We applied the same processing and augmentation for the modalities across all comparison methods. The preprocessed non-text modalities were tokenized using the tokenizer of the pre-trained vision encoder [28], following previous practices [9, 53].
A.2 Implementation details
We inserted prompts up to the fifth layer of each encoder, following [14]. To train modalities involving temporal information (video and audio), we duplicated the prompts for the temporal dimensions and concatenated them with the spatial features of each temporal dimension [29]. The modality features at time $t$, $V(e^{t}_{m,i};\hat{P}^{t}_{m})$, were average pooled along the temporal dimension before being fed into the projection head. To train the layer $g_{\beta^{t}}$, we computed and stored the mean and covariance of the features for each modality at the beginning of every time step. Specifically, for the $i$-th sample of modality $m$, $e^{t}_{m,i}$, the feature was calculated as $V(e^{t}_{m,i})$. For modality $m$, if the mean and covariance had already been stored from previous tasks, the mean was updated by combining the previous and new means based on their sample sizes, and the covariance was recomputed using the updated mean and new statistics. Finally, we trained the function $g_{\beta^{t}}$ using Equation (4) in the main paper.
Table A: Results for the reverse task order of the first (left) and second (right) sequences in the modality-agnostic class-incremental learning scenario. The AIA and F is measured after training on all tasks. Overall AIA and Overall F represent the average AIA and average F across all modalities.
Method Reverse Order of First Sequence (Random) Reverse Order of Second Sequence (Shift)
Audio ($\uparrow$) Depth ($\uparrow$) Video ($\uparrow$) Image ($\uparrow$)
Overall
AIA ($\uparrow$)
Overall
F ($\downarrow$)
Audio ($\uparrow$) Depth ($\uparrow$) Video ($\uparrow$) Image ($\uparrow$)
Overall
AIA ($\uparrow$)
Overall
F ($\downarrow$)
FT 27.18 26.65 9.13 6.33 17.32 8.75 12.24 16.89 20.13 38.98 22.06 17.93
EWC [16] 33.06 28.76 9.34 2.22 18.35 9.84 12.44 16.56 23.78 39.42 23.05 18.25
LwF [17] 26.12 22.26 13.88 8.06 17.58 11.35 12.90 17.45 24.67 39.91 23.73 19.16
WISE-FT [45] 28.48 28.13 8.71 4.57 17.47 9.63 14.16 17.42 24.67 38.57 23.71 16.81
L2P [44] 42.43 44.01 69.35 62.80 54.65 19.29 24.98 31.01 66.75 77.64 50.10 6.68
S-liPrompts [41] 44.03 45.94 74.03 61.14 56.29 20.01 24.85 31.18 71.65 75.10 50.70 12.04
AttriCLIP [40] 44.88 42.62 68.31 58.90 53.68 20.89 26.61 28.96 68.92 72.94 49.36 7.59
PGP [27] 41.20 49.40 79.63 79.30 62.39 17.05 26.08 33.56 62.13 79.82 50.40 11.37
COMM (Ours) 63.08 61.22 89.14 87.03 75.12 10.47 53.02 53.42 84.62 85.88 69.23 4.96
Table B: Results for three different class orders (mean $\pm$ std).
Method L2P [44] S-liPrompts [41] AttriCLIP [40] PGP [27] COMM
Image AIA ($\uparrow$) 79.3 $\pm$ 2.1 76.4 $\pm$ 0.6 77.9 $\pm$ 2.4 81.0 $\pm$ 1.8 85.9 $\pm$ 0.3
FAA ($\uparrow$) 70.9 $\pm$ 2.4 67.7 $\pm$ 1.8 66.6 $\pm$ 3.0 71.4 $\pm$ 2.6 79.6 $\pm$ 0.4
F ($\downarrow$) 8.8 $\pm$ 0.5 9.1 $\pm$ 1.4 11.9 $\pm$ 0.7 9.3 $\pm$ 1.1 6.3 $\pm$ 1.0
Video AIA ($\uparrow$) 80.6 $\pm$ 2.1 78.6 $\pm$ 3.9 80.4 $\pm$ 2.3 80.6 $\pm$ 2.5 90.1 $\pm$ 0.7
FAA ($\uparrow$) 69.8 $\pm$ 2.1 65.2 $\pm$ 3.0 72.0 $\pm$ 1.2 70.6 $\pm$ 0.6 82.9 $\pm$ 0.6
F ($\downarrow$) 11.5 $\pm$ 2.3 14.2 $\pm$ 1.3 8.9 $\pm$ 1.2 12.1 $\pm$ 1.8 7.7 $\pm$ 2.8
Depth AIA ($\uparrow$) 40.3 $\pm$ 3.8 44.5 $\pm$ 2.6 44.4 $\pm$ 4.8 46.2 $\pm$ 2.2 61.8 $\pm$ 2.7
FAA ($\uparrow$) 19.7 $\pm$ 4.4 23.5 $\pm$ 1.8 28.2 $\pm$ 4.8 29.1 $\pm$ 3.1 44.9 $\pm$ 1.0
F ($\downarrow$) 22.9 $\pm$ 1.4 24.5 $\pm$ 3.9 17.1 $\pm$ 1.1 18.1 $\pm$ 5.2 16.8 $\pm$ 3.8
Audio AIA ($\uparrow$) 46.7 $\pm$ 1.8 47.9 $\pm$ 4.5 46.4 $\pm$ 2.1 47.0 $\pm$ 0.9 65.7 $\pm$ 2.9
FAA ($\uparrow$) 19.5 $\pm$ 1.9 18.8 $\pm$ 2.9 20.5 $\pm$ 1.3 22.3 $\pm$ 3.0 47.9 $\pm$ 3.4
F ($\downarrow$) 28.9 $\pm$ 0.5 30.9 $\pm$ 6.1 27.5 $\pm$ 2.7 28.0 $\pm$ 2.1 18.9 $\pm$ 1.8
Overall AIA ($\uparrow$) 61.7 $\pm$ 2.4 61.9 $\pm$ 2.9 62.3 $\pm$ 2.9 63.7 $\pm$ 1.8 75.9 $\pm$ 1.6
FAA ($\uparrow$) 45.0 $\pm$ 2.7 43.8 $\pm$ 2.4 46.8 $\pm$ 2.5 48.3 $\pm$ 2.4 63.8 $\pm$ 1.3
F ($\downarrow$) 18.0 $\pm$ 1.1 19.7 $\pm$ 3.2 16.3 $\pm$ 1.4 16.8 $\pm$ 2.5 12.4 $\pm$ 1.7
B Additional Results
B.1 Results on different task orders
For the experiment conducted in the Table 1 in the main paper, the sequence consists of 20 tasks, where image, video, depth, and audio modalities were learned repetitively. For the sequence in Table 2, modality was trained for five time steps before shifting to next modality in the order of image, video, depth, and audio. To investigate the sensitivity of the compared methods with respect to different task orders in continual learning, we conducted an additional experiments using the reverse order of the tasks used in Tables 1 and 2.
The results are reported in Table A. The conventional continual learning methods, EWC, LwF, and WISE-FT, are highly sensitive to the order of tasks for both sequences. In contrast, the prompt-based methods, L2P, S-liPrompt, AttriCLIP, PGP, and COMM, show less sensitivity to the order of tasks compared to the conventional methods. The proposed method, COMM, demonstrates robustness to changes in task order, consistently outperforms other continual learning methods. Specifically, the proposed method shows a marginal overall AIA gap of 0.09% between the different task orders in the second sequence. However, other prompt-based continual learning methods, L2P, S-liPrompts, AttriCLIP, and PGP, show larger performance gaps of 3.64%, 2.31%, 5.67%, and 4.18%, respectively.
Figure A: t-SNE visualization on the modality embeddings extracted from the proposed method, COMM, with and without including self-regularization (SR) of prompts. Each color corresponds to an individual class.
B.2 Results on different class orders
We followed the class order provided by [37] for the video dataset. Similar to the practices in other continual learning benchmarks [37, 30, 40], we randomly selected the class order for the image [5], depth [35], and audio [25] datasets. Based on the determined class orders, the image and audio datasets were divided into five subsets consisting of 20 and 10 classes, respectively. The video dataset was divided into five subsets, with the first subset comprising 21 classes and the other four subsets containing 20 classes each. Similarly, the depth dataset was divided into five subsets, with the first four subsets comprising eight classes each and the remaining subset containing 12 classes. The class orders for the image, video, depth, and audio data used in the main paper are listed in Table C.
To validate the robustness of the proposed method under different class orders, we conducted additional experiments using three randomly shuffled class orders in the modality-specific scenario for the first sequence. Since regularization-based continual learning approaches exhibit unsatisfactory performance, we omit them from the comparison. Table B shows AIA, FAA, and Forgetting (F) results after training on all tasks. COMM consistently outperforms all other methods across all modalities in terms of AIA and FAA while exhibiting the lowest forgetting. Specifically, COMM outperforms the runner-up method, PGP, with performance margins of 12.2%, 15.5%, and 4.4% in AIA, FAA, and F, respectively. These results demonstrate the proposed COMM’s robustness to class order variations.
B.3 Impact of self-regularization
The proposed knowledge aggregation across time within a modality consolidates old and new knowledge by self-regulating the previously aggregated prompts and the current ones. This approach eliminates the need to store previous individual prompts $\{P^{1}_{m},\dots,P^{t-1}_{m}\}$ and $\{Q^{1}_{m},\dots,Q^{t-1}_{m}\}$, which provides computational efficiency compared to other prompt-based continual learning methods.
To clearly demonstrate the intrinsic maintainability of previous features by self-regularization (SR) after learning subsequent tasks, we visualize the embeddings obtained using the proposed method and the method without SR in Figure A. We compared the embeddings of tasks extracted right after training each task with those extracted after training all subsequent tasks. The results show that the embeddings remain almost unchanged even after learning new tasks compared to the previously extracted embeddings, when applying the presented SR. However, after learning subsequent tasks, embeddings from the method without SR become mixed, especially in the video and depth modalities.
Table C: Class orders (represented by their names) for each modality dataset.
Image Video
great white shark, drake, electric ray, mud turtle, sea lion, wallaby, conch, American coot, oystercatcher, barn spider, hognose snake, great grey owl, coucal, diamondback, prairie chicken, sea slug, black swan, black and gold garden spider, spiny lobster, leatherback turtle, black grouse, bittern, common newt, white stork, tiger shark, kite (bird of prey), tick, pelican, garden spider, green mamba, horned viper, sea anemone, snail, crayfish, vine snake, sulphur-crested cockatoo, hen, redshank, jellyfish, macaw, chambered nautilus, wolf spider, water ouzel, stingray, Dungeness crab, hammerhead, hermit crab, scorpion, crane, goldfinch, bustard, lorikeet, thunder snake, goldfish, tarantula, cock, agama, goose, ptarmigan, sea snake, flamingo, green snake, toucan, axolotl, chickadee, peacock, night snake, hornbill, tailed frog, common iguana, boa constrictor, red-backed sandpiper, Komodo dragon, spoonbill, black widow, bulbul, whiptail, rock crab, American alligator, flatworm, banded gecko, sidewinder, chiton, spotted salamander, magpie, bee eater, king snake, terrapin, hummingbird, wombat, garter snake, green lizard, tench, albatross, nematode, harvestman, bald eagle, indigo bunting, loggerhead, limpkin. surfing, salsa spin, playing tabla, pole vault, yo yo, mixing, apply eye makeup, volleyball spiking, long jump, field hockey penalty, rafting, hammering, push ups, handstand walking, blowing candles, golf swing, playing guitar, band marching, brushing teeth, playing flute, front crawl, archery, bench press, punch, horse riding, cutting in kitchen, billiards, cricket bowling, diving, haircut, boxing speed bag, horse race, still rings, high jump, biking, ice dancing, skijet, writing on board, hula hoop, throw discus, parallel bars, playing daf, wall pushups, swing, hammer throw, balance beam, typing, breast stroke, tennis swing, handstand pushups, sky diving, knitting, rope climbing, skiing, basketball dunk, baby crawling, rowing, nunchucks, table tennis shot, trampoline jumping, bowling, basketball, drumming, skate boarding, lunges, baseball pitch, kayaking, cliff diving, frisbee catch, rock climbing indoor, soccer juggling, shot put, playing piano, javelin throw, military parade, boxing punching bag, sumo wrestling, soccer penalty, pommel horse, walking with dog, clean and jerk, apply lipstick, playing violin, jumping jack, fencing, pull ups, juggling balls, body weight squats, pizza tossing, head massage, playing dhol, shaving beard, mopping floor, floor gymnastics, blow dry hair, jump rope, playing cello, playing sitar, tai chi, uneven bars, cricket shot.
Depth Audio
dancing room, computer room, coffee room, lobby, office dining, idk, storage room, bedroom, lecture theatre, library, corridor, reception room, playroom, classroom, cafeteria, gym, study space, indoor balcony, office, basement, living room, rest space, dinette, dining room, lab, furniture store, hotel room, printer room, bathroom, stairs, kitchen, reception, study, dining area, bookstore, exhibition, discussion area, home, mail room, laundromat, home office, recreation room, office kitchen, conference room. airplane, wind, clock tick, toilet flush, sneezing, rooster, sheep, pig, door wood creaks, hand saw, washing machine, crying baby, train, clock alarm, frog, rain, siren, crickets, snoring, helicopter, keyboard typing, cow, laughing, church bells, coughing, can opening, water drops, vacuum cleaner, insects, cat, door wood knock, thunderstorm, chainsaw, crackling fire, car horn, brushing teeth, footsteps, clapping, pouring water, mouse click, fireworks, breathing, sea waves, glass breaking, hen, chirping birds, crow, dog, drinking sipping, engine.